# Learning to Align Multi-Faceted Evaluation: A Unified and Robust Framework

### v2

@@ 1 Introduction
However, existing fine-tuned evaluators rely solely on text-based analysis with predefined evaluation criteria, leading to two key limitations Li et al. (2024b); Hu et al. (2024); Zhu et al. (2023); Kim et al. (2024b). First, evaluation based on predefined criteria can not fully capture the nuanced task requirements. For example, general criteria for writing, such as conciseness or logical structure, may not be sufficient for evaluating creative writing tasks that require an engaging plot. Moreover, it is challenging to effectively adapt predefined criteria to new and diverse instructions Li et al. (2024b). Second, LLM-based evaluators demonstrate significant instablity in evaluating adherence to complex instruction requirements, particularly objective criteria such as quantitative or structural constraints Zhou et al. (2023). For instance, they struggle to reliably assess basic textual attributes such as word counts, a common requirement in writing-related instructions Zhang and He (2024). These limitations also extend to the evaluation of formatting constraints.

### v3

@@ 1 Introduction
However, existing fine-tuned evaluators rely solely on text-based analysis with predefined evaluation criteria, leading to two key limitations Li et al. (2024b); Hu et al. (2024); Zhu et al. (2023); Kim et al. (2024b). First, evaluation based on predefined criteria can not fully capture the nuanced task requirements. For example, general criteria for writing, such as conciseness or logical structure, may not be sufficient for evaluating creative writing tasks that require an engaging plot. Moreover, it is challenging to effectively adapt predefined criteria to new and diverse instructions Li et al. (2024b). Second, LLM-based evaluators demonstrate significant instability in evaluating adherence to complex instruction requirements, particularly objective criteria such as quantitative or structural constraints Zhou et al. (2023). For instance, they struggle to reliably assess basic textual attributes such as word counts, a common requirement in writing-related instructions Zhang and He (2024). These limitations also extend to the evaluation of formatting constraints.
@@ 7 Conclusion
Acknowledgment
This work was supported by the Research Grants Council of Hong Kong (15207920, 15213323).
@@ Appendix B Prompt Templates
Appendix C Quality Analysis of Data Generation
Table 5 displays the success rate of each step in the filtering process.
Filtering Remaining rate
Check 1 (execute with the sample responses in the prompt) 100%
Check 2 (execute with another three sample responses) 99%
Reverse Validation 94%
Table 5: Filtering Process and Corresponding Remaining Rates
Overall, the high remaining rates across all steps indicate the effectiveness of our generation pipeline and the reliability of the final dataset. The reverse validation step, in particular, plays a crucial role in maintaining high-quality standards.
To enhance reproducibility, we have included representative examples of both accepted and rejected Python scripts in the supplementary material, along with the rationale for their inclusion or exclusion.
Appendix D Training Dataset Statistics
Table 6 presents a comprehensive summary of dataset statistics, including the distribution of task types, average length of analysis, and diversity of criteria. This provides clearer insight into the scope and representativeness of the training data.
Training Samples Quantity Average Length
Total 35144 111.5
- Text Analysis Samples 21060 81.7
- Python Script Samples 6322 247.9
- Evaluation Question Samples 7762 81.2
Table 6: Statistics of Training Samples by Category
Appendix E More Statistics of Type 2 Questions
Table 7 provides detailed statistics on Type 2 question coverage. Specifically, we categorize objective constraints into 8 types following Zhou et al. (2023) and report their proportions in the fine-tuning set. This addition helps illustrate how ARJudge generalizes across diverse constraint-based evaluations.
Constraint Type Count Percentage
Keywords 1,983 26.2%
Language 303 4.0%
Length Constraints 1,193 15.8%
Detectable Content 627 8.3%
Detectable Format 1,035 13.7%
Change Cases 374 4.9%
Start with / End with 1,741 23.0%
Punctuation 316 4.2%
Total 7572 100%
Table 7: Constraint Types with Corresponding Counts and Percentages
Appendix F Further Explanation of Positional Agreement Rate
While ARJudge shows a slightly lower Agr compared to Auto-J-13B, it is important to note that Auto-J-13B is nearly twice the size of our model and uses samples with swap responses to enhance positional agreement. When compared to models of similar scale—such as PandaLM-7B, Prometheus2-7B, and JudgeLM-7B—ARJudge achieves superior or comparable Average Agr scores as shown in Table 8.
Model Average Agr
GPT-4o 85.38
Claude-3.5-Sonnet 88.54
Deepseek-v3 85.30
Qwen2.5-7B 72.02
PandaLM-7B 76.65
Auto-J-13B 85.48
Prometheus2-7B 78.65
JudgeLM-7B 81.33
Themis-8B
ARJudge 80.85
Table 8: Average Agreement Scores Across Different Models
We believe the current level of positional agreement already demonstrates strong alignment. More importantly, ARJudge consistently outperforms all baselines across every evaluation benchmark, particularly on the more challenging LLMBar dataset (ARJudge Acc: 68.2, others < 42). This highlights the effectiveness of our multi-faceted analysis and refinement design, which contributes to more robust and accurate evaluations.
Appendix G Granular Ablation of Analysis Types
Table 9 presents a more fine-grained ablation study that evaluates a model variant with code-driven analysis disabled, in order to isolate the contribution of each analysis type. We do not evaluate a setting with only code-driven analysis, as most existing benchmarks are primarily designed for assessing content quality rather than objective constraints.
ARJudge MTBench (Acc) Auto-J (Acc) IFEval (Consistency)
Complete 78.3 78.5 85.6
Text Only 77.4 77.9 76.1
Table 9: Performance Comparison of ARJudge Methods across Different Benchmarks
The results demonstrate that code-driven analysis contributes meaningfully to performance, particularly on MTBench, Auto-J Eval, and IFEval. The variation in performance gains across these benchmarks reflects their differing focuses: MTBench and Auto-J Eval emphasize open-ended content evaluation, where textual coherence and reasoning dominate, while IFEval targets instruction-following with verifiable constraints, where code-based analysis plays a more central role.